OpenStudy (anonymous):

(x + 3)/4) * 3(x -6)/ 3(x + 3)

OpenStudy (anonymous):

cross cancel factors

OpenStudy (anonymous):

(x-6)/4

OpenStudy (anonymous):

\[\frac{ (x+3) }{ 4}*\frac{ 3(x-6) }{ 3(x+3) }\]

OpenStudy (anonymous):

(\[\frac{ (x-6) }{ 4 }\]

OpenStudy (anonymous):

can you show me how you get it @cebroski

OpenStudy (anonymous):

cross out the threes

OpenStudy (anonymous):

cross out the (x + 3)'s

OpenStudy (anonymous):

combine into one fraction

OpenStudy (jhannybean):

\[\large\frac{\cancel{(x+3)}}{4} \times \frac{\cancel{3}(x-6)}{\cancel{3(x+3)}}\]\[\large = \frac{x-6}{4}\]

OpenStudy (anonymous):

magick

OpenStudy (anonymous):

Okay Thanks everyone

OpenStudy (anonymous):

jhanny to the rescue. nice cross cancelling!!!

OpenStudy (jhannybean):

ty :)